Circular No. 8947

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA OPHIUCHI 2008 The following Japanese observers have reported independent discov-

eries on May 25 of an apparent tenth-magnitude nova: K. Nishiyama (Kurume, Fukuoka-ken) and F. Kabashima (Miyaki-cho, Saga-ken) on 60s frames taken with a 105-mm f/5.6 lens (communicated by S. Nakano, Sumoto, Japan); H. Nishimura (Kakegawa, Shizuoka-ken; via Nakano); and K. Haseda (Toyohashi, Aichi; via H. Yamaoka, Kyushu University). Nishimura and Haseda used 120-mm-f.l. camera lenses. Other confirming reports were received after word of the object was posted on the Central Bureau’s unconfirmed-objects webpage. CCD position measured for the apparent nova by Nishiyama and Kabashima (0.40-m f/9.8 reflector) on May 25.750 UT: α = 17^{h}39^{m}50^{s}.93, δ = −23^{o}50′00′′.9 (equinox 2000.0). Other measured position end figures: Nishimura, 51^{s}, 49′.1; Haseda, 47^{s}, 50′28′′ (resolution 14′′/pixel); S. Dvorak (near Orlando, FL, U.S.A., 0.25-m telescope on May 26.30), 50^{s}.90, 50′00′′.6; C. Jacques and E. Pimentel (Belo Horizonte, Brazil, 0.25-m reflector remotely near Mayhill, NM, U.S.A., on May 26.43), 50^{s}.92, 50′01′′.1; E. Guido and G. Sostero (Remanzacco, Italy, 0.25-m reflector, remotely near Mayhill, NM), 50^{s}.95, 50′00′′.9 (image posted at http://tinyurl.com/3v9cje). Magnitudes for the apparent nova (unfiltered CCD unless otherwise noted): 1985 (no date), [19.5 (Jacques and Pimentel, red Digitized Sky Survey); 1991 Aug. 11, [20: (Guido and Sostero, U.K. Schmidt Telescope red plate); 2008 Mar. 8.82, Apr. 4.77, and May 14.70, [11.5 (Haseda); Apr. 27.094, [16.0 (H. Mikuž, Crni Vrh Observatory; red magnitude); May 14.759, [12 (Nishimura); 19.752 and 20.694, [12.4 (Nishiyama and Kabashima); 20.745, [10 (Nishimura); 25.69, 10.3 (Haseda); 25.693 and 25.695, 10.2 (Nishiyama and Kabashima); 25.706, 10.2 (Nishimura); 25.750, 10.2 (Nishiyama and Kabashima); 26.30, V = 10.29, B = 11.72 (Dvorak); 26.31, B = 11.42, V = 10.45, R = 9.81 (Guido and Sostero); 26.43, 10.4 (Jacques and Pimentel). Visual magnitude estimates by A. Jones (Nelson, New Zealand): May 26.40, 10.2; 26.45, 10.1.

COMET 199P/SHOEMAKER Comet P/2008 G2 = 1994 J3 (cf. IAUC 8939) has been assigned the

number 199P (cf. MPC 62833).

2008 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green